Can you please assist me in this problem?
Find the axis of symmetry for y = x^2 - 9x +3.
Thanks
Neitra

Found 2 solutions by nerdybill, scott8148:
Answer by nerdybill(7384)

For ANY quadratic:
Ax^2 + Bx + C
the axis of symmetry is just:
x = -b/(2a)
(which is from the quadratic equation)
.
Therefore, for your problem:
Find the axis of symmetry for y = x^2 - 9x +3.
b is -9
a is 1
.
x = -(-9)/(2*1)
x = 9/2
It is the vertical line that crosses the x-axis at 9/2.

Answer by scott8148(6628)

the general equation is ___ x = -b / (2a)

in this case ___ x = -(-9) / 2 = 9/2